OpenStudy (anonymous):

what is d/dx(1/sinx)

OpenStudy (anonymous):

Gotta use the chain rule: \[\frac{d}{dx}[\frac{1}{f(x)}] = -\frac{1}{(f(x))^2}\cdot \frac{d}{dx}[f(x)]\] \[\implies \frac{d}{dx}[\frac{1}{sin(x)}] = -\frac{1}{sin^2(x)}\cdot cos(x) = -csc(x)cot(x)\]

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

By?

OpenStudy (anonymous):

ok,i hate calculas i just hate it,but thanks polpak and hasir for helping me

OpenStudy (anonymous):

Nothing to hate. Just have to practice and study.

OpenStudy (anonymous):

my SAT,i need to cry

OpenStudy (anonymous):

d/dx(1/sinx) take this rule: d/dx(a/u(x)) = -a.u'/u^2

OpenStudy (anonymous):

instead of making things more comlicated in memorizing